An extended source of light whose linear dimension is 0.3 mm is used in Young’s double slit experiment. Calculate the maximum separation between the slits for which interference pattern will be observable if the wavelength of light is 580 nm and separation between the source and the slits is 0.50 m.
1
Expert's answer
2020-02-07T10:29:26-0500

In YDSE condition for maximum is

"dsin\theta=\lambda" , where d is the separation between the slits

"d" "=\lambda/sin\theta"

"d=\lambda*" "D\/x"

"d=0.58*10^{-6}*0.5\/0.0003=0.97*10^{-3}m = 0.97 mm"